The beam supports the distributed load with wmax=3.6 kN/mwmax=3.6 kN/m as shown. The reactions at the supports AA and BB are vertical.

-- 1.5 m - 3 m — 1.5 m

Determine the resultant internal loadings acting on the cross section at point C.

Nc=?,Vc=?,Mc=?
Answer #1

Wmax = 3.6 kN/m wmáx?owa ANT amaz 3•6kN/m 1.5m t3m ti.sm/ RB For Equilibrium RA+RB= 0 x 6x316=10.8- EMB = 0, 62 RAX5 =%&*9x36For equation (1) Normal Force at -e x= 1.5 m Wa=Nc = 0.6 (6-1-5) = 2.7 kN/m Ne=2.7kN/m] AnyMoment at c. Me = 3.6x4.5-624.5x2.7%«4:5) = 16:2-9.1125 Me = 7.0875 kNm] Summary For triangular loading J DT TIL B Total load